OpenStudy (anonymous):

Calculate the concentration (in M) of 8.395 g of AgNO3 (molar mass = 169.9 g/mol) dissolved in 125 mL of water. Give your answer to three significant figures.

OpenStudy (anonymous):

So this is all about having the units cancel with the right conversions and the desired final units (M= mol/L) 8.395 (g) * 1/169.9 (mol/g) * 1/125 (1/ml) * 1000/1 (ml/L) = (8.395*1000)/(169.9*125) = 0.395 M (mol/L) Make sure you understand the conversions and how to invert the ratio if you want to flip the numerator/denominator. Make sure everything cancels and you get the desired units in the end. In this example, (g) cancels, (ml) cancels and you are left with (mol) on the numerator and (L) on the denominator. Please let me know if you have any other questions!